# Spectral upper bounds for the Grundy number of a graph

### 05 Jan 2024 v1

@@ 2 Grundy number and the largest root of the matching polynomial
For each positive integer $k$, a class of graphs denoted by $\mathbf{\mathcal{A}_{k}}$ was constructed in [15] which satisfies the following property. The Grundy number of any graph is at least $k$ if and only if $G$ contains an induced subgraph isomorphic to an element of $\mathbf{\mathcal{A}_{k}}$. Any element of $\mathbf{\mathcal{A}_{k}}$ is called $k$-atom. $K_{1}$ is the only 1-atom and $K_{2}$ is the only 2-atom. A graph $H$ is a $(k+1)$-atom if its vertices can be partitioned into an independent set $I$ and a $k$-atom $A_{k}$, such that each vertex of $A_{k}$ has exactly one edge incident to $I$ and each element from $I$ has at least one edge incident to $A_{k}$. We say that the seed of $A_{k}$ is a vertex that could have been the initial vertex during its recursive construction. For each $k$ there exists a unique tree that is a $k$-atom, denoted by $T_{k}$. We have that $|V(T_{k})|=2^{k-1}$ and that $T_{k}$ is the largest $k$-atom.
@@
where the sum runs over all $M\subseteq E$ which are matchings of the graph. Denote by $\phi_{G}(x)$ the characteristic polynomial of $A(G)$. The connection between the matching polynomial and the characteristic polynomial of a graph has been well understood and has found countless applications, including in the recent breakthrough [12]. A graph $G$ is a forest if and only if $\mu_{G}=\phi_{G}$ (see [8, Corollary 4.2]), but more generally, the roots of $\mu_{G}$ are all roots of the characteristic polynomial of the path tree of $G$, and, for any graph $G$ and vertex $u$, the following formula holds (see [8, Theorem 5]):
@@
For the result below, let $\mu_{1}(G)\geq\dots\geq\mu_{n}(G)$ denote the roots of the matching polynomial of a graph $G$. We will also use the fact that $\lambda_{1}(G)\geq\mu_{1}(G)$ for all graphs, and if $G$ is connected, equality holds if and only if $G$ is a tree ([8, Theorem 6]).
@@
The first inequality is [8, Theorem 6].
@@
Also, as a consequence of these facts and interlacing (see [5, Section 2.5]), the largest root of $\mu_{G}$ is strictly larger than any root of $\mu_{G\setminus u}$ for any vertex $u$. Thus, the second inequality in (3) holds because $A_{k}$ is a subgraph of $G$, with equality if and only if $G=A_{k}$.
@@ 3 Spectral upper bound depending on the size of the graph
Despite Theorem 4 implying tight limits for the Grundy number of trees $T_{k}$, it does not provide much information when the largest eigenvalue is not close to the square root of the maximum degree. Therefore, a question that arises is whether we can incorporate the number of vertices of the graph into the bound so as to balance this out.
This question has already been partially answered by considering the degeneracy of the graph. A graph $G$ is said to be $d$-degenerate if every subgraph has a vertex of degree at most $d$. The degeneracy of a graph is the smallest value of $d$ for which $G$ is $d$-degenerate. The Grundy number of $d$-degenerate graphs has been previously studied in [10], [1], and [6]. It has been proven that the Grundy number of $d$-degenerate graphs on $n$ vertices is bounded by:
@@ 4 Grundy number of random graphs
This already gives an indication of for which random graphs Theorem 4 derives a better bound than the maximum degree. Moreover, the same authors present the following result in [11, Corollary 1.2], concerning sparse random graphs. For any constant $c>0$, a.s.

### 23 Jan 2024 v2

@@ 2 Grundy number and the largest root of the matching polynomial
For each positive integer $k$, a class of graphs denoted by $\mathbf{\mathcal{A}_{k}}$ was introduced by Zeker [17], the class satisfies the following property. If the Grundy number of a graph $G$ is at least $k$, then $G$ contains a subgraph isomorphic to an element of $\mathbf{\mathcal{A}_{k}}$. Any element of $\mathbf{\mathcal{A}_{k}}$ is called $k$-atom. $K_{1}$ is the only 1-atom and $K_{2}$ is the only 2-atom. A graph $H$ is a $(k+1)$-atom if its vertices can be partitioned into an independent set $I$ and a $k$-atom $A_{k}$, such that each vertex of $A_{k}$ has exactly one edge incident to $I$ and each element from $I$ has at least one edge incident to $A_{k}$. In this paper, we use Zeker’s definition of $k$-atom, but slightly different definitions are present in the literature. For instance, Effantin et al. [8] define $H$ to be a $(k+1)$-atom if there is at least one edge from each vertex of $A_{k}$ to $I$. We state this difference because we do not use the complete characterization of Grundy numbers in terms of $k$-atoms from [17, Theorem 1], just the observation that if $\Gamma(G)\geq k$, then $G$ contains a $k$-atom as a subgraph.
We say that the seed of $A_{k}$ is a vertex that could have been the initial vertex during its recursive construction. For each $k$ there exists a unique tree that is a $k$-atom, denoted by $T_{k}$. We have that $|V(T_{k})|=2^{k-1}$ and that $T_{k}$ is the largest $k$-atom.
@@
where the sum runs over all $M\subseteq E$ which are matchings of the graph. Denote by $\phi_{G}(x)$ the characteristic polynomial of $A(G)$. The connection between the matching polynomial and the characteristic polynomial of a graph has been well understood and has found countless applications, including in the recent breakthrough [13]. A graph $G$ is a forest if and only if $\mu_{G}=\phi_{G}$ (see [9, Corollary 4.2]), but more generally, the roots of $\mu_{G}$ are all roots of the characteristic polynomial of the path tree of $G$, and, for any graph $G$ and vertex $u$, the following formula holds (see [9, Theorem 5]):
@@
For the result below, let $\mu_{1}(G)\geq\dots\geq\mu_{n}(G)$ denote the roots of the matching polynomial of a graph $G$. We will also use the fact that $\lambda_{1}(G)\geq\mu_{1}(G)$ for all graphs, and if $G$ is connected, equality holds if and only if $G$ is a tree ([9, Theorem 6]).
@@
The first inequality is [9, Theorem 6].
@@
Also, as a consequence of these facts and interlacing (see [5, Section 2.5]), the largest root of $\mu_{G}$ is strictly larger than any root of $\mu_{H}$ for any proper subgraph $H$. The reason for this is that the path tree $T(H,u)$ is a proper induced subgraph of the path tree $T(G,u)$, as the set of paths starting in $u$ in $H$ form a proper subset of the paths starting in $u$ in $G$. Thus, the second inequality in (3) holds because $A_{k}$ is a subgraph of $G$, with equality if and only if $G=A_{k}$.
@@ 3 Spectral upper bound depending on the size of the graph
Despite Theorem 4 implying tight limits for the Grundy number of trees $T_{k}$, it does not provide much information when the largest eigenvalue is not close to the square root of the maximum degree. Therefore, a question that arises is whether we can incorporate the size of the graph into the bound so as to balance this out.
Wu and Elphick [16] answer a question in this direction by finding an upper bound for the Grundy number in terms of the number of edges and the spectrum of the graph:
$\Gamma(G)\leq\frac{2|E(G)|}{\lambda_{1}(G)}.$
Moreover, the Grundy number can also be bounded only in terms of the largest eigenvalue and the number of vertices by using the degeneracy of the graph. A graph $G$ is said to be $d$-degenerate if every subgraph has a vertex of degree at most $d$. The degeneracy of a graph is the smallest value of $d$ for which $G$ is $d$-degenerate. The Grundy number of $d$-degenerate graphs has been previously studied in [11], [1], and [6]. It has been proven that the Grundy number of $d$-degenerate graphs on $n$ vertices is bounded by:
@@ 4 Grundy number of random graphs
This already gives an indication of for which random graphs Theorem 4 derives a better bound than the maximum degree. Moreover, the same authors present the following result in [12, Corollary 1.2], concerning sparse random graphs. For any constant $c>0$, a.s.